# Nonreciprocal transmission in hybrid atomic ensemble-optomechanical systems

### v2

@@ II.2 Dynamical equations
$\dot{a_{1}}$ $=-(i\Delta_{1}+\kappa_{1})a_{1}-ig_{1}a_{1}(b+b^{\dagger})-iJ_{1}a_{2}-iJ_{2}d+\mathcal{E}_{1}+\mathcal{E}_{p1}e^{-i\delta t},$ (14a)
@@
$\mathcal{E}_{p1}^{out}+\mathcal{E}^{in}_{p1}e^{-iyt}=\sqrt{\kappa_{1}}\delta\tilde{a}_{1},$ (26)
$\mathcal{E}_{p2}^{out}+\mathcal{E}^{in}_{p2}e^{-iyt}=\sqrt{\kappa_{2}}\delta\tilde{a}_{2},$
where $\mathcal{E}_{p1,p2}^{in}=\mathcal{E}_{p1,p2}/\sqrt{\kappa_{j}}$. For simplicity, we assume that the output fields also take the form:
@@
$\mathcal{E}^{out}_{p1+}=\sqrt{\kappa_{1}}\delta\tilde{a}_{1+}-\frac{\mathcal{E}_{p1}}{\sqrt{\kappa_{1}}},$ $\mathcal{E}_{p1-}^{out}=0,$ (28)
$\mathcal{E}^{out}_{p2+}=\sqrt{\kappa_{2}}\delta\tilde{a}_{2+}-\frac{\mathcal{E}_{p2}}{\sqrt{\kappa_{2}}},$ $\mathcal{E}_{p2-}^{out}=0.$
@@ III.1 Optical nonreciprocal transmission
Figure 2: Density plot of transmission amplitudes ($\mathcal{T}_{12}$ and $\mathcal{T}_{21}$) versus relative phase difference $\theta$ between effective optomechanical coupling strength and the phase $\phi$. We have used $G_{1}/\gamma=0.5$, $G_{2}=G_{1}$, $f/\gamma=10$, $\kappa_{1}/\gamma=1$, $\kappa_{2}=\kappa_{1}$, $J_{1}/\gamma=0.5$, $\omega_{m}/\gamma=10$, $\Delta^{\prime}_{1}=\Delta^{\prime}_{2}=\Delta_{\rm en}=\omega_{m}$, $|J_{2}|/\gamma=0.01$, and $J_{3}/\gamma=4.476e^{i\frac{\pi}{2}}$.
@@ III.2 Influence of coupling strength and detuning
Figure 4: Transmission amplitudes $\mathcal{T}_{12}$ and $\mathcal{T}_{21}$ as functions of detuning $y/\gamma$ for different coupling constants. (a) $|J_{2}|/\gamma=0.01$, (b) $|J_{2}|/\gamma=0.1$, (c) $|J_{2}|/\gamma=0.3$, (d) $|J_{2}|/\gamma=0.5$, (e) $J_{3}/\gamma=0.476e^{i\frac{\pi}{2}}$, (f) $J_{3}/\gamma=1.476e^{i\frac{\pi}{2}}$, (g) $J_{3}/\gamma=2.476e^{i\frac{\pi}{2}}$, and (h) $J_{3}/\gamma=4.476e^{i\frac{\pi}{2}}$. The other parameters are the same as in Figure 2, except for $\theta=\phi=\frac{\pi}{2}$.
@@ III.4 Influence of atomic ensemble decay and mechanical damping
Figure 9: Transmission amplitudes $\mathcal{T}_{12}$ (black solid–line) and $\mathcal{T}_{21}$ (blue dotted–line) as a function of detuning $y/\gamma$ for different values of $J_{3}$. Parameters used are the same as for Figure 2 except for $\theta=\phi=\frac{\pi}{2}$ and complex coupling $J_{3}$ (a) $J_{3}/\gamma=4.676e^{i\frac{\pi}{2}}$, (b) $J_{3}/\gamma=4.276e^{i\frac{\pi}{2}}$
@@ Acknowledgments
Author Contributions: K.B.E. and J.-X. conceptualized the work and carry out the simulations and analysis. A., P.D., and S.G.N.E. participated in all the discussions and provided useful suggestions to the final version of the manuscript. K.S.N. and A.-H.A.-A. supervised the work. All authors participated equally in the discussions and the preparation of the final manuscript.
@@
$\mathcal{D}$ $=-2J_{1}|J_{2}|J_{3}G_{2}\cos(\theta-\phi)-2i|J_{2}|J_{3}\kappa_{2}G_{1}\cos\phi$ (31)
@@
$+y^{2}\mathcal{D}_{7}+\kappa_{1}\kappa_{2}\mathcal{D}_{8}-i\gamma fy\mathcal{D}_{9}+y^{4},$
@@
For simplicity, we consider the case where $\theta=\phi=\frac{\pi}{2}$, and the expression of $\mathcal{D}$ becomes $\mathcal{D}=J_{3}^{2}(R_{5}+R_{2}^{\prime})+R_{4}+\frac{R_{6}}{J_{3}^{2}}(J_{1}J_{3}^{2}+J_{1}\gamma f-G_{1}G_{2}f)^{2}+R_{3}$, where $R_{1}=\frac{1}{\sqrt{\kappa_{1}\kappa_{2}}}$, $R_{2}=\frac{-2G_{1}G_{2}f}{R_{1}}$, $R^{\prime}_{2}=J_{1}^{2}+\kappa_{1}\kappa_{2}$, $R_{3}=G^{2}_{2}f\kappa_{1}+G_{1}^{2}f\kappa_{2}+J_{1}^{2}\gamma f+\kappa_{1}\kappa_{2}\gamma f$, $R_{4}=-2J_{1}f(J_{1}\gamma-G_{1}G_{2})$, $R_{5}=-2J_{1}^{2}$, $R_{6}=\frac{G_{2}^{2}+\gamma\kappa_{2}}{G_{2}^{2}}$, $R_{7}=R_{5}+R_{2}^{\prime}+R_{6}J_{1}^{2}$, $R_{8}=R_{4}-R_{2}+2R_{6}J_{1}^{2}\gamma f-2R_{6}J_{1}G_{1}G_{2}f+R_{3}$, and $R_{9}=R_{6}J_{1}^{2}\gamma^{2}f^{2}-2R_{6}J_{1}\gamma f^{2}G_{1}G_{2}+R_{6}G_{1}^{2}G_{2}^{2}f^{2}$.

### v3

@@ II.2 Dynamical equations
$\dot{a_{1}}$ $=-(i\Delta_{1}+\kappa_{1})a_{1}-ig_{1}a_{1}(b+b^{\dagger})-iJ_{1}a_{2}-iJ_{2}d$
$\quad+\mathcal{E}_{1}+\mathcal{E}_{p1}e^{-i\delta t},$ (14a)
@@
$\mathcal{E}_{p1}^{out}+\mathcal{E}^{in}_{p1}e^{-iyt}=\sqrt{2\kappa_{1}}\delta\tilde{a}_{1},$ (26)
$\mathcal{E}_{p2}^{out}+\mathcal{E}^{in}_{p2}e^{-iyt}=\sqrt{2\kappa_{2}}\delta\tilde{a}_{2},$
where $\mathcal{E}_{p1,p2}^{in}=\mathcal{E}_{p1,p2}/\sqrt{2\kappa_{j}}$. For simplicity, we assume that the output fields also take the form:
@@
$\mathcal{E}^{out}_{p1+}=\sqrt{2\kappa_{1}}\delta\tilde{a}_{1+}-\frac{\mathcal{E}_{p1}}{\sqrt{2\kappa_{1}}},$ $\mathcal{E}_{p1-}^{out}=0,$ (28)
$\mathcal{E}^{out}_{p2+}=\sqrt{2\kappa_{2}}\delta\tilde{a}_{2+}-\frac{\mathcal{E}_{p2}}{\sqrt{2\kappa_{2}}},$ $\mathcal{E}_{p2-}^{out}=0.$
@@ III.1 Optical nonreciprocal transmission
Figure 2: Density plot of transmission amplitudes ($\mathcal{T}_{12}$ and $\mathcal{T}_{21}$) versus relative phase difference $\theta$ between effective optomechanical coupling strength and the phase $\phi$. We have used $G_{1}/\gamma=0.5$, $G_{2}=G_{1}$, $f/\gamma=10$, $\kappa_{1}/\gamma=2$, $\kappa_{2}=\kappa_{1}$, $J_{1}/\gamma=0.5$, $\omega_{m}/\gamma=10$, $\Delta^{\prime}_{1}=\Delta^{\prime}_{2}=\Delta_{\rm en}=\omega_{m}$, $|J_{2}|/\gamma=0.01$, and $J_{3}/\gamma=4.346e^{i\frac{\pi}{2}}$.
@@ III.2 Influence of coupling strength and detuning
Figure 4: Transmission amplitudes $\mathcal{T}_{12}$ and $\mathcal{T}_{21}$ as functions of detuning $y/\gamma$ for different coupling constants. (a) $|J_{2}|/\gamma=0.01$, (b) $|J_{2}|/\gamma=0.1$, (c) $|J_{2}|/\gamma=0.3$, (d) $|J_{2}|/\gamma=0.5$, (e) $J_{3}/\gamma=0.346e^{i\frac{\pi}{2}}$, (f) $J_{3}/\gamma=1.346e^{i\frac{\pi}{2}}$, (g) $J_{3}/\gamma=2.346e^{i\frac{\pi}{2}}$, and (h) $J_{3}/\gamma=4.346e^{i\frac{\pi}{2}}$. The other parameters are the same as in Figure 2, except for $\theta=\phi=\frac{\pi}{2}$.
@@ III.4 Influence of atomic ensemble decay and mechanical damping
Figure 9: Transmission amplitudes $\mathcal{T}_{12}$ (black solid–line) and $\mathcal{T}_{21}$ (blue dotted–line) as a function of detuning $y/\gamma$ for different values of $J_{3}$. Parameters used are the same as for Figure 2 except for $\theta=\phi=\frac{\pi}{2}$ and complex coupling $J_{3}$ (a) $J_{3}/\gamma=4.646e^{i\frac{\pi}{2}}$, (b) $J_{3}/\gamma=4.126e^{i\frac{\pi}{2}}$
@@ Acknowledgments
Author Contributions: E.K.B. and J.-X. conceptualized the work and carry out the simulations and analysis. A., P.D., and S.G.N.E. participated in all the discussions and provided useful suggestions to the final version of the manuscript. K.S.N. and A.-H.A.-A. supervised the work. All authors participated equally in the discussions and the preparation of the final manuscript.
@@
$\mathcal{D}$ $=2(-2J_{1}|J_{2}|J_{3}G_{2}\cos(\theta-\phi)-2i|J_{2}|J_{3}\kappa_{2}G_{1}\cos\phi$ (31)
@@
$+y^{2}\mathcal{D}_{7}+\kappa_{1}\kappa_{2}\mathcal{D}_{8}-i\gamma fy\mathcal{D}_{9}+y^{4}),$
@@
For simplicity, we consider the case where $\theta=\phi=\frac{\pi}{2}$, and the expression of $\mathcal{D}$ becomes $\mathcal{D}=2(J_{3}^{2}(R_{5}+R_{2}^{\prime})+R_{4}+\frac{R_{6}}{J_{3}^{2}}(J_{1}J_{3}^{2}+J_{1}\gamma f-G_{1}G_{2}f)^{2}+R_{3})$, where $R_{1}=\frac{1}{\sqrt{\kappa_{1}\kappa_{2}}}$, $R_{2}=\frac{-2G_{1}G_{2}f}{R_{1}}$, $R^{\prime}_{2}=J_{1}^{2}+\kappa_{1}\kappa_{2}$, $R_{3}=G^{2}_{2}f\kappa_{1}+G_{1}^{2}f\kappa_{2}+J_{1}^{2}\gamma f+\kappa_{1}\kappa_{2}\gamma f$, $R_{4}=-2J_{1}f(J_{1}\gamma-G_{1}G_{2})$, $R_{5}=-2J_{1}^{2}$, $R_{6}=\frac{G_{2}^{2}+\gamma\kappa_{2}}{G_{2}^{2}}$, $R_{7}=R_{5}+R_{2}^{\prime}+R_{6}J_{1}^{2}$, $R_{8}=R_{4}-R_{2}+2R_{6}J_{1}^{2}\gamma f-2R_{6}J_{1}G_{1}G_{2}f+R_{3}$, and $R_{9}=R_{6}J_{1}^{2}\gamma^{2}f^{2}-2R_{6}J_{1}\gamma f^{2}G_{1}G_{2}+R_{6}G_{1}^{2}G_{2}^{2}f^{2}$.